A line up for tickets to a local concert had an average (mean) waiting time of 20 minutes with a  standard deviation of 4 minutes. 

a. What percentage of the people in line waited for more than 20 minutes? 




1
Expert's answer
2022-03-29T07:17:51-0400

"P\left( X>20 \right) =P\left( \frac{X-20}{4}>\frac{20-20}{4} \right) =P\left( Z>0 \right) =1-\varPhi \left( 0 \right) =0.5"